# Langevin equations and a geometric integration scheme for the overdamped limit of homogeneous rotational Brownian motion

### v2

@@ III Geometric derivation of the overdamped Langevin equation
with $\tau_{\text{R}}^{-1}:=(d-1)D_{\text{R}}$ for motion in $d$-dimensional space ($d=2,3$). The transformation is straightforward [52], but a bit tedious due to the vector cross product on the r.h.s. of Eq. 7 and will be given below, after an alternative derivation which is based on geometric considerations. We will also show that, due to the specific rules of Itō calculus, the additional term $-\tau_{\text{R}}^{-1}\boldsymbol{u}(t)$ does not generate a drift, but is necessary to preserve the norm $|\boldsymbol{u}(t)|$ and to yield the orientational autocorrelation function correctly. Furthermore, the solution $\boldsymbol{u}(t)$ is non-anticipating with respect to the noise $\boldsymbol{\omega}(t)$, i.e., the (Itō) increments of the noise, $\int_{t}^{t+\Delta t}\boldsymbol{\omega}(s)\mathrm{d}s$ are independent of $\boldsymbol{u}(s)$ for any $t\geqslant s$ and $\Delta t>0$. (This central statement has no simple analogue for the Stratonovich integral.) As a consequence, the Itō–Langevin equation (9) is amenable to the straightforward Euler–Maruyama scheme for numerical integration [81]. Below, we propose a geometric integration scheme that satisfies the constraint $|\boldsymbol{u}(t)|=1$ exactly for arbitrary $\Delta t$.
III.1 Geometric derivation of the Itō form
In the following, we adopt a geometric perspective to derive Eq. 9. For Brownian motion of the vector $\boldsymbol{u}(t)$ on the three-dimensional unit sphere, the evolution of $\boldsymbol{u}(t)$ emerges from the repeated action of infinitesimal random rotations, which are the analogue of the random displacements of planar Brownian motion. Thus, we can write:
@@ III.1 Geometric derivation of the Itō form
$\boldsymbol{\omega}\times\boldsymbol{v}=(\boldsymbol{\omega}\cdot\boldsymbol{J})\boldsymbol{v}$ (12)
@@
Expanding the exponential in Eq. 10 up to the order of $[\boldsymbol{\omega}(t)\mathrm{d}t]^{2}$ yields, according to the rules of Itō calculus,
@@
Equation 13 is the same as Eq. 9, which is seen by subtracting $\boldsymbol{u}(t)$ from both sides of Eq. 13 and dividing it by the differential $\mathrm{d}t$.
@@ III.2 Discussion of the drift term
Before, we discuss the term $-\tau_{\text{R}}^{-1}\boldsymbol{u}(t)$ on the r.h.s. of Eq. 9, which appears to describe a drift. However, the term is normal to the constraining manifold and hence cannot generate a drift. Instead, it is needed to satisfy the constraint, i.e., to preserve the normalisation $|\boldsymbol{u}(t)|=1$ under the dynamics of Eq. 9. In order to prove this fact, we switch to the notation of stochastic (Itō) differentials and write $\mathrm{d}\boldsymbol{u}(t)=\dot{\boldsymbol{u}}(t)\mathrm{d}t$; the expression $\boldsymbol{\omega}(t)\mathrm{d}t$ is the differential of a scaled (and shifted  [^3] The argument given holds also in the presence of $\boldsymbol{T}_{\text{ext}}(t)$ in Eq. 6.) Wiener process. Then, Eq. 6 implies $\omega_{i}(t)\mathrm{d}t\,\omega_{j}(t)\mathrm{d}t=2D_{\text{R}}\delta_{ij}\mathrm{d}t$ for the Cartesian components $i$ and $j$ of $\boldsymbol{\omega}(t)$. Employing Itō’s formula, substituting $\mathrm{d}\boldsymbol{u}(t)$ by means of Eq. 9, and omitting the time arguments for brevity, one finds:
@@
=(9)${{=}}-2\tau_{\text{R}}^{-1}\boldsymbol{u}\cdot\boldsymbol{u}\mathrm{d}t+(\boldsymbol{\omega}\mathrm{d}t\times\boldsymbol{u})\cdot(\boldsymbol{\omega}\mathrm{d}t\times\boldsymbol{u})$ )
@@ III.3 Orientational autocorrelation function
The importance of the Itō term in Eq. 9 becomes clear again when computing the autocorrelation function of the orientation, $C_{1}(t):=\langle\boldsymbol{u}(t)\cdot\boldsymbol{u}(0)\rangle$, which is a simple exercise in the absence of an external torque $\boldsymbol{T}_{\text{ext}}(t)=0$. Multiplication of Eq. 9 by $\boldsymbol{u}(0)$, integrating over time, and averaging yields
@@
Recalling that $\boldsymbol{\omega}(s)$ is a white noise process [Eq. 6] and independent of $\boldsymbol{u}(s)$ and $\boldsymbol{u}(0)$, the first term on the r.h.s. contains a properly formed Itō integral, which is zero on average. (The Stratonovich integral does not share this property.) Here, we have used that $\langle\omega(t)\rangle=\boldsymbol{T}_{\text{ext}}(t)=0$. The remaining integral equation has the expected solution
@@ III.4 Equivalence of the Itō and Stratonovich forms
It remains to show the mathematical equivalence of Eqs. 7 and 9. Generally, given a diffusion process $\boldsymbol{X}(t)$ in $\mathbb{R}^{d}$, which is driven by a standard Wiener process $\boldsymbol{W}(t)$ in $\mathbb{R}^{n}$ scaled by an $d\times n$ matrix-valued coefficient function $\sigma(\boldsymbol{X}(t))$, the corresponding Stratonovich integral, denoted by $\hbox{}\circ\mathrm{d}\boldsymbol{W}(t)$ as usual, is related to the Itō integral via [79, 51]
@@
for some time $T>0$ and $\partial_{k}$ denoting the partial derivative w.r.t. to the component $X_{k}$. Omitting the integral signs in Eq. 18 (formally, taking a derivative w.r.t. $T$) yields the differential form of this relation in the common short-hand notation:
@@
In order to apply it to Eq. 7, one identifies $\boldsymbol{X}(t)$ with $\boldsymbol{u}(t)$ and $\mathrm{d}\boldsymbol{W}(t)$ with $(2D_{\text{R}})^{-1/2}\boldsymbol{\omega}(t)\mathrm{d}t$. The noise strength $\sigma(\cdot)$ is $3\times 3$ matrix-valued, and it is convenient to use the above matrix form of the vector product: $\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}\boldsymbol{u}\cdot\boldsymbol{J}$, which has the derivatives $\partial_{k}\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{k}$. It follows that
$[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{ij}\circ\omega_{j}(t)\,\mathrm{d}t$ $=[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{ij}\,\omega_{j}(t)\,\mathrm{d}t$
@@
$=\boldsymbol{\omega}(t)\,\mathrm{d}t\times\boldsymbol{u}(t)-2D_{\text{R}}\boldsymbol{u}(t)\,\mathrm{d}t\,;$ (20)
in the last step, we have simplified the last term as follows:
@@
$\delta_{ij}$ is the Kronecker symbol. After division of Eq. 20 by $\mathrm{d}t$, it is clear that, for $\tau_{\text{R}}^{-1}=2D_{\text{R}}$, the r.h.s. of the Stratonovich SDE (7) corresponds to the r.h.s. of the Itō SDE (9) as claimed.
@@ III.5 Two-dimensional case
The geometric derivation of Eq. 9 has a straightforward generalisation to $d=2$ dimensions. We note that the angular velocity $\boldsymbol{\omega}(t)$, being an axial vector, has $n:=d(d-1)/2$ components since it can be represented as a skew-symmetric $d\times d$ matrix. In particular for $d=2$, it has only a single component $\omega_{z}$ (which may be chosen as the $z$-component in a three-dimensional embedding of the motion). Correspondingly, the Lie algebra $\mathfrak{so}(2)$ is generated by a single matrix, e.g., $J_{z}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. With this, the cross product in Eq. 12 is to be understood as
@@
for two-dimensional vectors $\boldsymbol{v}$. The reasoning from Eqs. 10 to 13 remains unchanged except for the last step, where one has to replace $-\sum_{i}(J_{i})^{2}=2\cdot\mathds{1}$ by $-(J_{z})^{2}=\mathds{1}$. Hence, Eq. 9 applies also in $d=2$ dimensions, but with $\tau_{\text{R}}^{-1}=D_{\text{R}}$.
For the proof of the equivalence of the Itō and Stratonovich forms (Section III.4), the following changes apply: The noise strength $\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{z}\boldsymbol{u}$ is a $2\times 1$ matrix with derivative $\partial_{k}\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{z}\boldsymbol{e}_{k}$; here, $\boldsymbol{e}_{k}$ are the Cartesian unit vectors, $(\boldsymbol{e}_{k})_{j}=\delta_{kj}$. Second, the calculation in Eq. 21 is replaced by
@@ IV.1 Algorithm
Above, we have shown that Eq. 10 is equivalent to Eq. 9. The latter is a standard Itō stochastic differential equation and may be integrated numerically with the Euler–Maruyama scheme or some higher-order scheme [81, 51], introducing some finite integration time step $\Delta t$. However, these schemes would satisfy the constraint $|\boldsymbol{u}(t)|=1$ only asymptotically, for $\Delta t\to 0$.
Alternatively, Eq. 10 suggests to implement a geometric integration scheme for Brownian motion on the unit sphere as a sequence of rotations with finite random angles, which reads for a single integration step:
@@
$\boldsymbol{u}(t+\Delta t)=\cos(\vartheta)\boldsymbol{u}-\sin(\vartheta)\,\boldsymbol{u}\times\boldsymbol{n}+[1-\cos(\vartheta)](\boldsymbol{u}\cdot\boldsymbol{n})\boldsymbol{n}\,.$ (25)
We note that, nowadays, the trigonometric functions are evaluated by dedicated special function units on typical computing hardware, either CPUs or GPUs, and the appearance of such functions in the algorithm is not a performance issue anymore.
It remains to specify the statistics of $\Delta\boldsymbol{\Omega}(t)$. Comparison of Eq. 24 and Eq. 10 yields for small time steps that
@@
With this choice, time integration of Eq. 6 suggests that $\Delta\boldsymbol{\Omega}(t)$ is a Gaussian vector with independent components. Clearly, the rotation of $\boldsymbol{u}$ about its own axis has no effect, which also follows from setting $\boldsymbol{n}=\boldsymbol{u}$ in Eq. 25. Thus, it is sufficient to choose $\Delta\boldsymbol{\Omega}$ in the plane perpendicular to $\boldsymbol{u}$:
@@
Table 1: Algorithm for the geometric integration scheme [Eq. 24] using Gaussian distributed random rotations [Eq. 28].
In summary, each time step of the geometric integration scheme [Eqs. 24, 25, 26 and 28] consists of the algorithm listed in Table 1. In step (iv), we have simplified Eq. 25, noting that $\boldsymbol{u}\perp\boldsymbol{n}$. Optionally, after step (iv), one may normalise the unit vector $\boldsymbol{u}(t+\Delta t)$ to avoid a possible drift of $|\boldsymbol{u}|$ after a large number of integration steps, which can result from round-off errors due to limited floating-point precision. We emphasise again that the time discretisation in this scheme preserves the normalisation exactly.

### v3

@@ II.4 Itō form of the overdamped equation
with $\tau_{\text{R}}^{-1}:=(d-1)D_{\text{R}}$ for motion in $d$-dimensional space ($d=2,3$). The transformation is straightforward [52], but a bit tedious due to the vector cross product on the r.h.s. of Eq.˜7 and will be given below, after an alternative construction based on geometric considerations. We will also show that, due to the specific rules of Itō calculus, the additional term $\tau_{\text{R}}^{-1}\boldsymbol{u}(t)$ does not generate a drift, but is necessary to preserve the norm $|\boldsymbol{u}(t)|$ and to yield the orientational autocorrelation function correctly. Furthermore, the solution $\boldsymbol{u}(t)$ is non-anticipating with respect to the noise $\boldsymbol{\omega}(t)$, i.e., the (Itō) increments of the noise, $\int_{t}^{t+\Delta t}\boldsymbol{\omega}(s)\mathrm{d}s$ are independent of $\boldsymbol{u}(s)$ for any $t\geqslant s$ and $\Delta t>0$. (This central statement has no simple analogue for the Stratonovich integral.) As a consequence, the Itō–Langevin equation (9) is amenable to the straightforward Euler–Maruyama scheme for numerical integration [63]. Below, we propose a geometric integration scheme that satisfies the constraint $|\boldsymbol{u}(t)|=1$ exactly for arbitrary $\Delta t$.
III Geometric construction of the overdamped Langevin equation
The discussion in Section˜II is given from the angle of a bottom-up approach, starting from a Hamiltonian many-body system (Brownian particle plus solvent particles) and systematically integrating out fast degrees of freedom. Here, we adopt a geometric perspective and follow a top-down approach by postulating properties of (idealised) Brownian motion on the unit sphere, similarly as it is established for Brownian motion in flat, Euclidean spaces. We then show that this geometric construction [cf. Eq.˜10] is equivalent to the accepted SDE for overdamped rotational Brownian motion, Eq.˜7.
III.1 Geometric construction and derivation of the Itō form
In the following, we adopt a geometric perspective to obtain Eq.˜9. The Wiener process $\boldsymbol{W}(t)$, i.e., free and idealised Brownian motion, in a flat, Euclidean space can be constructed as a Lévy process with independent, stationary, and Gaussian distributed increments (or: displacements) $\boldsymbol{W}(t^{\prime})-\boldsymbol{W}(t)$ for $t^{\prime}>t$. Analogously, Brownian motion of the orientation vector $\boldsymbol{u}(t)$ on the (three-dimensional) unit sphere is a Lévy process with independent, stationary, and isotropic increments such that $\boldsymbol{u}(t^{\prime})\cdot\boldsymbol{u}(t)$ samples a certain distribution [cf. Eq.˜38] determined by the diffusion equation on the sphere [cf. Eq.˜53]. Here, the dot product enters via the usual metric on the unit sphere: the Euclidean distance is replaced by the length of the great circle arc between $\boldsymbol{u}(t^{\prime})$ and $\boldsymbol{u}(t)$, which is $\arccos(\boldsymbol{u}(t^{\prime})\cdot\boldsymbol{u}(t))$. One concludes that the evolution of $\boldsymbol{u}(t)$ emerges from the repeated action of infinitesimal random rotations and we write:
@@ III.1 Geometric construction and derivation of the Itō form
$\boldsymbol{\omega}\times\boldsymbol{v}=(\boldsymbol{\omega}\cdot\boldsymbol{J})\boldsymbol{v}=(\boldsymbol{v}\cdot\boldsymbol{J})^{\top}\boldsymbol{\omega}$ (12)
@@
Expanding the exponential in Eq.˜10 up to the order of $\boldsymbol{\omega}(t)\mathrm{d}t]^{2}$ yields, according to the rules of Itō calculus,
@@
Equation˜13 is the same as Eq.˜9, which is seen by subtracting $\boldsymbol{u}(t)$ from both sides of Eq.˜13 and dividing it by the differential $\mathrm{d}t$.
@@ III.2 Discussion of the drift term
Before, we discuss the term $\tau_{\text{R}}^{-1}\boldsymbol{u}(t)$ on the r.h.s. of Eq.˜9, which appears to describe a drift. However, the term is normal to the constraining manifold and hence cannot generate a drift. Instead, it is needed to satisfy the constraint, i.e., to preserve the normalization $|\boldsymbol{u}(t)|=1$ under the dynamics of Eq.˜9. In order to prove this fact, we switch to the notation of stochastic (Itō) differentials and write $\mathrm{d}\boldsymbol{u}(t)=\dot{\boldsymbol{u}}(t)\mathrm{d}t$; the expression $\boldsymbol{\omega}(t)\mathrm{d}t$ is the differential of a scaled (and shifted  [^3] The argument given holds also in the presence of $\boldsymbol{T}_{\text{ext}}(t)$ in Eq. 6.) Wiener process. Then, Eq.˜6 implies $\omega_{i}(t)\mathrm{d}t\,\omega_{j}(t)\mathrm{d}t=2D_{\text{R}}\delta_{ij}\mathrm{d}t$ for the Cartesian components $i$ and $j$ of $\boldsymbol{\omega}(t)$. Employing Itō’s formula, substituting $\mathrm{d}\boldsymbol{u}(t)$ by means of Eq.˜9, and omitting the time arguments for brevity, one finds:
@@
[…]
@@
In Section˜VI, we derive the Fokker–Planck equation corresponding to Eq.˜9. From a different angle, it also proves that the term $\tau_{\text{R}}^{-1}\boldsymbol{u}(t)$ does not contribute to the probability flux. (In fact, the absence of this term would entail a spurious drift.) Moreover, for $\boldsymbol{T}_{\text{ext}}=0$, the equilibrium distribution of the orientation $\boldsymbol{u}$ is obtained to be uniform on the unit sphere, as required for free rotational diffusion.
@@ III.3 Orientational autocorrelation function
The importance of the Itō term in Eq.˜9 becomes clear again when computing the autocorrelation function of the orientation, $C_{1}(t):=\langle\boldsymbol{u}(t)\cdot\boldsymbol{u}(0)\rangle$, which is a simple exercise in the absence of an external torque $\boldsymbol{T}_{\text{ext}}(t)=0$. Multiplication of Eq.˜9 by $\boldsymbol{u}(0)$, integrating over time, and averaging yields
@@
Recalling that $\boldsymbol{\omega}(s)$ is a white noise process [Eq.˜6] and independent of $\boldsymbol{u}(s)$ and $\boldsymbol{u}(0)$, the first term on the r.h.s. contains a properly formed Itō integral, which is zero on average. (The Stratonovich integral does not share this property.) Here, we have used that $\langle\omega(t)\rangle=\boldsymbol{T}_{\text{ext}}(t)=0$. The remaining integral equation has the expected solution
@@ III.4 Equivalence of the Itō and Stratonovich forms
It remains to show the mathematical equivalence of Eqs.˜7 and 9. Generally, given a diffusion process $\boldsymbol{X}(t)$ in $\mathbb{R}^{d}$, which is driven by a standard Wiener process $\boldsymbol{W}(t)$ in $\mathbb{R}^{n}$ scaled by an $d\times n$ matrix-valued coefficient function $\sigma(\boldsymbol{X}(t))$, the corresponding Stratonovich integral, denoted by $\hbox{}\circ\mathrm{d}\boldsymbol{W}(t)$ as usual, is related to the Itō integral via [80, 51]
@@
for some time $T>0$ and $\partial_{k}$ denoting the partial derivative w.r.t. to the component $X_{k}$; the summation convention is applied to repeated indices. Omitting the integral signs in Eq.˜18 (formally, taking a derivative w.r.t. $T$) yields the differential form of this relation in the common short-hand notation:
@@
In order to apply it to Eq.˜7, one identifies $\boldsymbol{X}(t)$ with $\boldsymbol{u}(t)$ and $\mathrm{d}\boldsymbol{W}(t)$ with $(2D_{\text{R}})^{-1/2}\delta\boldsymbol{\omega}(t)\,\mathrm{d}t$. The noise strength $\sigma(\cdot)$ is $3\times 3$ matrix-valued: $\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}\,\boldsymbol{u}\cdot\boldsymbol{J}$, which has the derivatives $\partial_{k}\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{k}$. This is seen by rewriting the cross product in matrix form [Eq.˜12], $(\delta\boldsymbol{\omega}\times\boldsymbol{u})\,\mathrm{d}t=-(\boldsymbol{u}\cdot\boldsymbol{J})\,\delta\boldsymbol{\omega}\,\mathrm{d}t$. Therefore,
$[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{ij}\circ\delta\omega_{j}(t)\,\mathrm{d}t\hskip-50.00008pt$
$=[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{ij}\,\delta\omega_{j}(t)\,\mathrm{d}t$
@@
$=\delta\boldsymbol{\omega}(t)\,\mathrm{d}t\times\boldsymbol{u}(t)-2D_{\text{R}}\boldsymbol{u}(t)\,\mathrm{d}t\,.$ (20)
For the last step, we have simplified as follows:
@@
$\delta_{ij}$ is the Kronecker symbol. After division of Eq.˜20 by $\mathrm{d}t$, it is clear that, for $\tau_{\text{R}}^{-1}=2D_{\text{R}}$, the r.h.s. of the Stratonovich SDE (7) corresponds to the r.h.s. of the Itō SDE (9) as claimed. We note that the external torque contributes to the systematic drift and does not enter $\sigma(\boldsymbol{u})$ and, thus, not the transformation (19) of the stochastic integral; in particular, $\delta\boldsymbol{\omega}(t)$ in Eq.˜20 may be replaced by $\boldsymbol{\omega}(t)$.
@@ III.5 Two-dimensional case
The geometric construction of Eq.˜9 has a straightforward generalization to $d=2$ dimensions. We note that the angular velocity $\boldsymbol{\omega}(t)$, being an axial vector, has n:=d(d1)/2n:=d(d-1)/2 components since it can be represented as a skew-symmetric $d\times d$ matrix. In particular for $d=2$, it has only a single component $\omega_{z}$ (which may be chosen as the $z$-component in a three-dimensional embedding of the motion). Correspondingly, the Lie algebra $\mathfrak{so}(2)$ is generated by a single matrix, e.g., $J_{z}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. With this, the cross product in Eq.˜12 is to be understood as
@@
for two-dimensional vectors $\boldsymbol{v}$. The reasoning from Eqs. 10 to 13 remains unchanged except for the last step, where one has to replace $\sum_{i}(J_{i})^{2}=2\cdot\mathds{1}$ by $(J_{z})^{2}=\mathds{1}$. Hence, Eq.˜9 applies also in $d=2$ dimensions, but with $\tau_{\text{R}}^{-1}=D_{\text{R}}$.
For the proof of the equivalence of the Itō and Stratonovich forms (Section˜III.4), the following changes apply: The noise strength $\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{z}\boldsymbol{u}$ is a $2\times 1$ matrix with derivative $\partial_{k}\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{z}\boldsymbol{e}_{k}$; here, $\boldsymbol{e}_{k}$ are the Cartesian unit vectors, $(\boldsymbol{e}_{k})_{j}=\delta_{kj}$. Second, the calculation in Eq.˜21 is replaced by
@@ IV.1 Algorithm
Above, we have shown that Eq.˜10 is equivalent to Eq.˜9. The latter is a standard Itō SDE and may be integrated numerically with the Euler–Maruyama scheme or some higher-order scheme [63, 51], introducing some finite integration time step $\Delta t$. However, these schemes would satisfy the constraint $|\boldsymbol{u}(t)|=1$ only asymptotically, for $\Delta t\to 0$.
Alternatively, Eq.˜10 suggests to implement a geometric integration scheme for Brownian motion on the unit sphere as a sequence of rotations with finite random angles, which reads for a single integration step:
@@
$\boldsymbol{u}(t+\Delta t)=\cos(\vartheta)\boldsymbol{u}-\sin(\vartheta)\,\boldsymbol{u}\times\boldsymbol{n}\\ +[1-\cos(\vartheta)](\boldsymbol{u}\cdot\boldsymbol{n})\boldsymbol{n}\,.$ (25)
Nowadays, the trigonometric functions are evaluated by dedicated special function units on typical computing hardware, either CPUs or GPUs, and the appearance of such functions in the algorithm is not a performance issue anymore.
It remains to specify the statistics of $\Delta\boldsymbol{\Omega}(t)$. Comparison of Eq.˜24 and Eq.˜10 yields for small time steps that
@@
With this choice, time integration of Eq.˜6 suggests that $\Delta\boldsymbol{\Omega}(t)$ is a Gaussian vector with independent components. Clearly, the rotation of $\boldsymbol{u}$ about its own axis has no effect, which also follows from setting $\boldsymbol{n}=\boldsymbol{u}$ in Eq.˜25. Thus, it is sufficient to choose $\Delta\boldsymbol{\Omega}$ in the plane perpendicular to $\boldsymbol{u}$:
@@
Table 1: Algorithm for the geometric integration scheme [Eq.˜24] using Gaussian distributed random rotations [Eq.˜28].
In summary, each time step of the geometric integration scheme [Eqs.˜24, 25, 26 and 28] consists of the algorithm listed in Table˜1. In step (iv), we have simplified Eq.˜25, noting that $\boldsymbol{u}\perp\boldsymbol{n}$. Optionally, after step (iv), one may normalize the unit vector $\boldsymbol{u}(t+\Delta t)$ to avoid a possible drift of $|\boldsymbol{u}|$ after a large number of integration steps, which can result from round-off errors due to limited floating-point precision. We emphasize again that the time discretization in this scheme preserves the normalization exactly.